OpenStudy (anonymous):

Clear fractions or decimals, solve, and check. -5/6 + x= -1/2 -2/3

OpenStudy (anonymous):

0

OpenStudy (anonymous):

-5/6 + x= -1/2 -2/3 -5/6+x=-3/6-4/6 -5/6+x=-(3+4)/6 -5/6+x=-7/6 x=-1/3

OpenStudy (anonymous):

i want to you clear it? Does me understand right?

OpenStudy (anonymous):

clear what

OpenStudy (anonymous):

_5/6+x= -(1/2+2/3).....LCM of 2 and 3 =6.....so -5/6+x=-(3+2/6) -5/6+x= -(5/6).....take -5/6 to the other side, the sign changes to 5/6..so we have x=-5/6+5/6 =0 x=0........

OpenStudy (anonymous):

dose -5/6 + x=-5/(6+x) or just (-5/6 )+ x

OpenStudy (anonymous):

it's (-5/6)+ x the question was not written well by the person that posted it

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

it don't have () that in it

OpenStudy (anonymous):

even if you do it like that, you will still end up getting zero

OpenStudy (anonymous):

no the answer is -1/3

OpenStudy (anonymous):

ok so the final answer is -1/3